# Dynamic Importance in Diffusion U-Net for Enhanced Image Synthesis

### v1

@@ I Introduction
Previous studies have revealed the roles of the attention blocks in the U-Net. It can be empirically summarized that, high-resolution blocks primarily focus on detail extraction, while mid-low-resolution blocks correspond to layout structuring and semantic understanding [5]. Subsequent works on plug-and-play attention features show that, the U-Net also attends to features of different gratuities at variant denoising time steps [6, 7, 8]. Recently, FreeU [9] attempted to analyze the functionality of the attention blocks, showing that the backbone features and the skip connections of the U-Net contribute to information of different frequencies. Based on this finding, a re-weighting scheme is proposed to enhance the generation quality. However, they overlook the dynamic shifts in block roles during the denoising process, which hinders their further exploration.
@@ II-A U-Net Mechanisms in Diffusion Models
Recently, there has been growing interest in the interpretability of diffusion models, especially the functionality of U-Net. For instance, [10] proposes a hypothesis regarding the specific role of each layer within U-Net. Additionally, some research has explored U-Net’s mechanism from the frequency domain. FreeU [9] examines the component variation of different frequencies during the denoising process, pointing out that the U-Net backbone primarily contributes to denoising, while skip connections introduce high-frequency features into the decoder. Similarly, [11] found DMs are inclined to generate high-frequency features, and learn to recover components of varying frequencies at different time steps.
@@ III-A Preliminary
Given a clean sample $\boldsymbol{x}_{0}$ and a variance schedule $\{\overline{\alpha}_{i}\}_{i=1}^{n}$, the deterministic reverse step of DDIM [2] is
@@
where $\boldsymbol{\epsilon}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$ is the true noise, and $\hat{\boldsymbol{\epsilon}}_{t}=\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x}_{t},t)$ represents the noise predicted by the U-Net parameterized by $\theta$ at time step $t$.
@@
$\text{SNR}(\boldsymbol{x}_{t})=||\boldsymbol{x}_{0}||^{2}/\text{Var}(\Delta\boldsymbol{\epsilon}_{t})=||\boldsymbol{x}_{0}||^{2}/\text{Var}(\boldsymbol{\epsilon}-\hat{\boldsymbol{\epsilon}}_{t}),$ (2)
where $||\boldsymbol{x}_{0}||^{2}$ represents the power of the true signal $\boldsymbol{x}_{0}$, and $\Delta\boldsymbol{\epsilon}_{t}=\boldsymbol{\epsilon}-\hat{\boldsymbol{\epsilon}}_{t}$ denotes the error between the true noise $\boldsymbol{\epsilon}$ and the predicted noise $\hat{\boldsymbol{\epsilon}}_{t}$.
@@
$\boldsymbol{y}_{i}=\boldsymbol{f}_{i}(\boldsymbol{x}_{0})+\boldsymbol{g}_{i}(\boldsymbol{\epsilon})+\boldsymbol{n}_{i},$ (3)
@@
  • $\boldsymbol{g}_{i}(\boldsymbol{\epsilon})$: The feature components related to the noise $\boldsymbol{\epsilon}$.
@@
$\begin{split}&\text{Var}(\Delta\hat{\boldsymbol{\epsilon}})\approx\textstyle\sum_{i}A_{i}^{2}(w_{i}-1)^{2}\text{Var}(\boldsymbol{g}_{i}(\boldsymbol{\epsilon}))\\ &~{}~{}~{}~{}+\textstyle\sum_{i}A_{i}^{2}w_{i}^{2}\text{Var}(\boldsymbol{f}_{i}(\boldsymbol{x}_{0}))+\textstyle\sum_{i}A_{i}^{2}w_{i}^{2}\text{Var}(\boldsymbol{n}_{i}),\end{split}$ (4)
@@ III-B Re-weighting the Outputs of Transformer Blocks
Intuitively, applying a weight $w>1.0$ amplifies the effect of the attention mechanism [15] within the Transformer block, whereas applying a weight $w<1.0$ attenuates it. More rigorously, in accordance with Prop. 1, we aim to reduce $\text{Var}(\Delta\hat{\boldsymbol{\epsilon}})$ via re-weighting, thus enhancing the SNR.
However, we empirically observed that assigning arbitrary weights greater that 1.0 to any block does not always lead to performance enhancement (see Tab. I). In some cases, it can even yield worse performance compared to the vanilla weighting schedule, i.e., $w_{i}=1$ for all blocks.
We note that, this arises because the importance of blocks dynamically shifts throughout the denoising process (see Sec. IV-B). Statically assigning fixed weights to certain blocks may misweight the contributions of components in Eq. 4, thus increasing $\text{Var}(\Delta\hat{\boldsymbol{\epsilon}})$, and consequently reducing the SNR. This led us to uncover and quantify the dynamic importance of each Transformer block thoughout the denoising process.
@@ III-D Quantify Block Importance via the Voting Mechanism
in which $r$ is the number of runs.
@@ IV-D Enhanced Image Synthesis
Firstly, we fix $high=1.1$ for the weight range, and investigate the impact of varying $l\in[0.95,1.05]$. The variation of aesthetic scores with respect to $l$ is plotted in Fig. 3. The results indicate that our re-weighting schedule consistently outperforms the vanilla one in both the training and test set, demonstrating the robustness and generalization of our method across different categories of prompts.

### v2

@@ I Introduction
Previous studies have revealed the roles of the attention blocks in the U-Net. It can be empirically summarized that, high-resolution blocks primarily focus on detail extraction, while mid-low-resolution blocks correspond to layout structuring and semantic understanding [5]. Subsequent works on plug-and-play attention features show that, the U-Net also attends to features of different gratuities at variant denoising time steps [6, 7, 8]. Recently, FreeU [9] attempted to analyze the functionality of the attention blocks, showing that the backbone features and the skip connections of the U-Net contribute to information of different frequencies. Based on this finding, a re-weighting scheme is proposed to enhance the generation quality. However, they overlook the dynamic shifts in block roles during the denoising process, which hinders their further exploitation.
@@ II-A U-Net Mechanisms in Diffusion Models
Recently, there has been growing interest in the interpretability of diffusion models, especially the functionality of U-Net. For instance, [10] proposes a hypothesis regarding the specific role of each layer within U-Net. Researchers have also explored U-Net’s mechanism from the frequency domain. FreeU [9] examines the component variation of different frequencies during the denoising process, pointing out that the U-Net backbone primarily contributes to denoising, while skip connections introduce high-frequency features into the decoder. Similarly, [11] found DMs are inclined to avoid generating high-frequency features, and learn to recover components of varying frequencies at different time steps.
@@ III-A Preliminary
Given a clean sample $\boldsymbol{x}_{0}$ and a variance schedule $\{\overline{\alpha}_{t}\}_{t=0}^{T}$, the deterministic reverse step of DDIM [2] is
@@
where $T$ is the amount of training steps of DMs, and $\hat{\boldsymbol{\epsilon}}_{t}=\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x}_{t},t)$ represents the noise predicted by the U-Net parameterized by $\theta$ at time step $t$.
@@
$\text{SNR}(\boldsymbol{x}_{t})=||\boldsymbol{x}_{0}||^{2}/\text{Var}(\Delta\boldsymbol{\epsilon}_{t})=||\boldsymbol{x}_{0}||^{2}/\text{Var}(\boldsymbol{\epsilon}_{t}-\hat{\boldsymbol{\epsilon}}_{t}),$ (2)
where $||\boldsymbol{x}_{0}||^{2}$ represents the power of the true signal $\boldsymbol{x}_{0}$, and $\Delta\boldsymbol{\epsilon}_{t}=\boldsymbol{\epsilon}_{t}-\hat{\boldsymbol{\epsilon}}_{t}$ denotes the error between the true noise $\boldsymbol{\epsilon}_{t}$ and the predicted noise $\hat{\boldsymbol{\epsilon}}_{t}$.
@@
$\boldsymbol{y}_{i}=\boldsymbol{f}_{i}(\boldsymbol{x}_{0})+\boldsymbol{g}_{i}(\boldsymbol{\epsilon}_{t})+\boldsymbol{n}_{i},$ (3)
@@
  • $\boldsymbol{g}_{i}(\boldsymbol{\epsilon}_{t})$: The feature components related to the noise $\boldsymbol{\epsilon}_{t}$.
@@
$\begin{split}&\text{Var}(\Delta\hat{\boldsymbol{\epsilon}}_{t})\approx\textstyle\sum_{i}A_{i}^{2}(w_{i}-1)^{2}\text{Var}(\boldsymbol{g}_{i}(\boldsymbol{\epsilon}_{t}))\\ &~{}~{}~{}~{}+\textstyle\sum_{i}A_{i}^{2}w_{i}^{2}\text{Var}(\boldsymbol{f}_{i}(\boldsymbol{x}_{0}))+\textstyle\sum_{i}A_{i}^{2}w_{i}^{2}\text{Var}(\boldsymbol{n}_{i}),\end{split}$ (4)
@@ III-B Re-weighting the Outputs of Transformer Blocks
Intuitively, applying a weight $w>1.0$ amplifies the effect of the attention mechanism [15] within the Transformer block, whereas applying a weight $w<1.0$ attenuates it. More rigorously, in accordance with Prop. 1, we aim to reduce $\text{Var}(\Delta\hat{\boldsymbol{\epsilon}}_{t})$ via re-weighting, thus enhancing the SNR.
However, we empirically observed that assigning arbitrary weights greater that 1.0 to any block does not always lead to performance enhancement (see Tab. I). In some cases, it can even yield worse performance compared to the vanilla weighting schedule, i.e., $w_{i}=1.0$ for all blocks.
We note that, this arises because the importance of blocks dynamically shifts throughout the denoising process (see Sec. IV-B). Statically assigning fixed weights to certain blocks may misweight the contributions of components in Eq. 4, thus increasing $\text{Var}(\Delta\hat{\boldsymbol{\epsilon}}_{t})$, and consequently reducing the SNR. This led us to uncover and quantify the dynamic importance of each Transformer block thoughout the denoising process.
@@ III-D Quantify Block Importance via the Voting Mechanism
in which $r$ is the number of runs, $vs_{i}^{(t)}$ is the voting score.
@@ IV-D Enhanced Image Synthesis
Firstly, we fix $high=1.1$ for the weight range, and investigate the impact of varying $low\in[0.95,1.05]$. The variation of aesthetic scores with respect to $low$ is plotted in Fig. 3. The results indicate that our re-weighting schedule consistently outperforms the vanilla one in both the training and test set, demonstrating the robustness and generalization of our method across different categories of prompts.